OpenStudy (anonymous):

systems of equations y=4 third x plus 3 y= negative x minus 4

OpenStudy (anonymous):

so y=4/3x+3 y=-x-4 soooo, let's find the x! in the second equation for example: x=-y-4 ok? Sostitute this x in the first equation => y=4/3*(-y-4)+3 resolve => y=-4/3y-16/3+3 => 3/3y+4/3y=-16/3+9/3 sooo: 7/3y=-5 => y=-15/7 tadaaaaa you find y! now sostitute this y to the second equation: -15/7=-x-4 => x=-28/7+15/7 => x=-13/7 !!! finish ;)

OpenStudy (anonymous):

i don't follow

OpenStudy (anonymous):

i don't see why you swapped out the -x for -y

OpenStudy (anonymous):

?i don't undertand, sorry, but I'm not english XD

OpenStudy (anonymous):

and thats not even the right answer

OpenStudy (anonymous):

x = -3

OpenStudy (anonymous):

probably I do wrong calculus, but the main idea is this

OpenStudy (anonymous):

i have the answers just want to know how to do the problem

OpenStudy (anonymous):

and its not even calculus its algebra 1

OpenStudy (anonymous):

i intend calculate for caluclus XD

OpenStudy (anonymous):

so, if you're ready we try to make another time this sistem